What is the rate of reaction when [N2O5]=0.0250[N2O5]=0.0250 ?

Express the rate in molarity per second to three significant figures.

1
Expert's answer
2022-03-16T10:51:03-0400

[N2O5] = 0.0250 M;

For the complete definition of rate of reaction one must know the course of the reaction and the reaction rate constant (K).

If the reaction is the zero order reaction then: v = K(M/s) * [N2O5]0 = K * 1 = K (M/s);

If the reaction is the first order reaction then: v = K(1/s) * [N2O5]1 = K * [N2O5] = K * 0.025 (M/s);

If the reaction is the second order reaction then: v = K(1/(M*s)) * [N2O5]2 = K * [N2O5] * [N2O5] = K * 0.025 * 0.025 = K * 0.001 (M/s).